EXOTIC METAL MOLECULES IN OXYGEN-RICH ENVELOPES: DETECTION OF AlOH (X^{1}Σ^{+}) IN VY CANIS MAJORIS

E. D. TENENBAUM and \underline{L. M. ZIURYS}, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

A new interstellar molecule, AlOH, has been detected toward the envelope of VY Canis Majoris, an oxygen-rich red supergiant. Three rotational transitions of AlOH were observed using the facilities of the Arizona Radio Observatory (ARO). The J = 9 → 8 and J = 7 → 6 lines at 1 mm were measured with the ARO Submillimeter Telescope (SMT), while the J = 5 → 4 transition at 2 mm was observed with the ARO 12 m antenna on Kitt Peak. The AlOH spectra exhibit quite narrow line widths, indicating that the emission arises from within the dust acceleration zone of the central circumstellar outflow. From a radiative transfer analysis, the abundance of AlOH relative to H_{2} was found to be 1x10^{−7} for a source size of 0.26′′ or 22 R_{∗}. AlOH is likely formed just beyond the photosphere via thermodynamic equilibrium chemistry, and then disappears due to dust condensation. The AlOH/AlO abundance ratio found in VY CMa is ∼17. LTE calculations predict the monohydroxides should be the major carriers of Al, Ca, and Mg in O-rich envelopes, as opposed to the oxides or halides.